Your cable company offer two plans: basic service with one movie channel for $35 per month or basic service with two movie channel $45 per month. What is the charge for basic service and charge for each movie channel
Answer by jorel1380(3719)

f+a=35
f+2a=45
2a-a=a=10
f=25
The basic service is $25--Each movie channel is $10 more..